OpenStudy (anonymous):

are: x^3 and x^2*|x|, Linear independent?

OpenStudy (anonymous):

\[(x^{3}) and( x^{2}|x|)\]

OpenStudy (anonymous):

no (i don't think) because you can write \[x^2|x| = x^3 \]or \[x^2|x|=-x^3\]

OpenStudy (anonymous):

yea, that is also what i assume, so I have to prove that there are no a and b so that(except the trivial case a=0 and b=0): \[a*(x3)+b*(x2|x|)=0\]

OpenStudy (anonymous):

and a and b must be in \[\mathbb{R}\]

OpenStudy (anonymous):

what space are you working in?

OpenStudy (anonymous):

\[[-1,1]\rightarrow \mathbb{R}\]

OpenStudy (anonymous):

because in fact it is possible to have \[ax^3+bx^2|x|=0\] say x = -3, a = 1, b = -1

OpenStudy (anonymous):

oh so space of continuous functions from \[[-1,1]->R\]

OpenStudy (anonymous):

those functions are in vectorspace of all functions, and to be shown is that they are linear independent if they are. I assumption was also all the time that they are independent But how would one go and prove it

OpenStudy (anonymous):

yea [−1,1]−>R

OpenStudy (anonymous):

i guess you have to work in cases. \[ x^2|x| = \left\{ \begin{array}{lr} x^3 & : x \geq 0\\ -x^3 & : x < 0 \end{array} \right.\]

OpenStudy (anonymous):

and the equality \[ax^3+bx^2|x|=0\] means \[ax^3+bx^2|x|=0 \forall x\in [-1,1]\]

OpenStudy (anonymous):

my thought was: \[x ^{2}|x|\] this is always positive, and because \[x^{3}\] can be also negative there is no any single a that could make \[a∗(x3)+b∗(x2|x|)=0\] true

OpenStudy (anonymous):

a should be dependable about x (if <0 or if >0)

OpenStudy (anonymous):

so since clearly you have positive values of x in [-1,1] you get \[ax^3+bx^2|x|=0\] mean in particular that for all \[x \in [0,1]\] \[ax^3+bx^3=0\]

OpenStudy (anonymous):

AND \[ax^3-bx^3=0\] \[\forall x \in [-1,0)\]

OpenStudy (anonymous):

both have to be true, since it is true \[\forall x \in [-1,1]\]

OpenStudy (anonymous):

yes true and therefore I can't define any value a and b that would generally make it hold hold.

OpenStudy (anonymous):

the first equality says \[(a+b)x^3=0 \forall x \in [0,1]\] therefore \[a=-b\]

OpenStudy (anonymous):

the second says \[(a-b)x^3=0 \forall x \in [-1,0)\] therefore \[a=b\]

OpenStudy (anonymous):

and the only way that \[a=b\] and \[a=-b\] is if b = 0, hence so is a

OpenStudy (anonymous):

But I understand that I should be after all be able to define a and b so that they would make it hold for every case -1 =< x = <1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

otherwise they are linear independent

OpenStudy (anonymous):

like the case here?

OpenStudy (anonymous):

the logic goes like this: assume \[ax^3+bx^2|x|=0 \forall x \in [-1,1]\]

OpenStudy (anonymous):

oh damn i have just tried to show dependence and i showed independence!

OpenStudy (anonymous):

i have showed that if \[ax^3+bx^2|x|=0 \forall x \in [-1,1]\] then \[a=b=0\]

OpenStudy (anonymous):

so maybe my first assumption was wrong!

OpenStudy (anonymous):

but that would be the case here i think since we cant define any specific a and b that would hold for -1 < x < 1

OpenStudy (anonymous):

since there is a need to seperate 2 cases for x negative and positive

OpenStudy (anonymous):

i always think of it as this: can you write one as a linear combination of the other. the formal definition says the same thing. so the question is can we write \[x^2|x|\] as some linear combination of \[x^3\] and i guess we cannot.

OpenStudy (anonymous):

since one is a piecewise function and the other is not

OpenStudy (anonymous):

i guess they are independent after all!

OpenStudy (anonymous):

yes I think so too

OpenStudy (anonymous):

whew! learn something new every day. mainly that i should not jump to a conclusion. since i thing the logic was sound and we have shown that if \[ax^3+bx^2|x|=0 \forall x \in [-1,1]\] then \[a=b=0\]

OpenStudy (anonymous):

pretty sure actually, we would need a function in a place of a or b to make it work, but since it must be from \[R\] there is no way to make it work

OpenStudy (anonymous):

then done. good work.

OpenStudy (anonymous):

ok, thanks for your help